OpenStudy (toxicsugar22):

how do u solve 2/x+3-6/x-3=12/x^2-9

OpenStudy (anonymous):

\[\frac{2}{x+3}-\frac{6}{x-3}=\frac{12}{x^2-9}\]

OpenStudy (anonymous):

multiply both sides by \[x^2-9\]

OpenStudy (anonymous):

on the right you will get 12. on the left you will get \[2(x-3)-6(x+3)\] because \[x^2-9=(x+3)(x-3)\]

OpenStudy (anonymous):

so the factors will cancel. now solve \[2(x-3)-6(x+3)=12\] \[2x-6-6x-18=12\] \[-4x-24=12\] \[-4x=12+24=36\] \[x=\frac{36}{-4}=-9\]

OpenStudy (anonymous):

let me check it to makes sure i didn't make a mistake

OpenStudy (anonymous):

looks right

OpenStudy (toxicsugar22):

so the anwer is -9

OpenStudy (anonymous):

yeah

OpenStudy (toxicsugar22):

yah i got that

OpenStudy (anonymous):

that is what i got but i still want to check with the original equation to see that i did not make a careless algebra mistake

OpenStudy (toxicsugar22):

can u help me with this one

OpenStudy (toxicsugar22):

nm

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

btw it is right i checked. get \[\frac{1}{6}\] on both left and right

OpenStudy (toxicsugar22):

so the aswere is -9 right

OpenStudy (toxicsugar22):

so its the anwer -9